Two bodies were thrown simultaneously from the same point: one, straight up, and the other, at an angle of θ = 60° to the horizontal. The initial velocity of each body is equal to v0 = 25 m/s. Neglecting the air drag, find the distance between the bodies t = 1.70 s later.
1
Expert's answer
2020-10-15T03:13:11-0400

"v_0 = 25m\/s"

"t = 1.70s"

Let the body thrown straight up be 1, and the other one be 2

Note that;

"a_y = -g = -\,10m/s^2"

"ax = 0\,m/s^2"


For body 1

"\theta"1 = "90\text{\textdegree}" (i.e vertical)

"\begin{aligned}

U_{y1} = v_0sinθ \quad &; \quad U_{x1} = v_0cosθ\\
U_{y1} = 25sin90 \enspace &; \quad U_{x1} = 25cos90\\
U_{y1} = 25m/s \quad & ;\quad U_{x1} = 0m/s

\end{aligned}"


"\begin{alignedat}{2}

&x_1 = U_{x1}t\, &+& \,0.5a_xt^2\\
&x_1 = 0(1.70)\, &+&\, 0.5(0)(1.70)^2\\
&x_1 = 0 + 0\\
∴& x_1 = 0m

\end{alignedat}"


"\begin{alignedat}{2}

&y_1 = U_{y1}t \,&+& \,0.5a_yt^2\\
&y_1 = (25)(1.70) \,&+&\, 0.5(-10)(1.70)^2\\
&y_1 = 42.5\, &-&\, 14.45\\
∴& y_1 = 28.05m

\end{alignedat}"



For body 2

"\theta"2 = "60\text{\textdegree}"

"\begin{aligned}

U_{y2} = v_0sinθ \quad \,\,\, &; \quad U_{x2} = v_0cosθ\\
U_{y2}= 25sin60 \quad & ; \quad U_{x2} = 25cos60\\
U_{y2}= 21.65m/s\,\, & ; \quad U_{x2} = 12.5m/s

\end{aligned}"


"\begin{alignedat}{2}

&x_2 = U_{x2}t \, &+&\, 0.5a_xt^2\\
&x_2 = 12.5(1.70) \,&+& \,0.5(0)(1.70)^2\\
&x_2 = 21.25 &+&\, 0\\
∴ &x_2 = 21.25m

\end{alignedat}"


"\begin{alignedat}{2}

&y_2 = U_{y2}t \,&+&\, 0.5a_yt^2\\
&y_2 = (21.65)(1.70)\, &+&\, 0.5(-10)(1.70)^2\\
&y_2 = 36.81 &-& \,14.45\\
∴&y_2= 22.36m

\end{alignedat}"


from, distance between 2 points = "\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}"


distance between the 2 bodies = "\sqrt{(21.25-0)^2+(22.36-28.05)^2}"


= "\sqrt{(21.25)^2+(-5.69)^2}"


= "\sqrt{451.56+32.38}"


= "\sqrt{483.94}"


= "22m"


"\therefore" the distance between the two bodies is 22m